### v2

# Wigner distribution, Wigner entropy, and Anomalous Transport of a Generalized Aubry-André model

Abstract
In this paper, we study a generalized Aubry-André model with tunable quasidisordered potentials. The model has an invariable mobility edge that separates the extended states from the localized states. At the mobility edge, the wave function presents critical characteristics, which can be verified by finite-size scaling analysis. Our numerical investigations demonstrate that the extended, critical, and localized states can be effectively distinguished via their phase space representation, specially the Wigner distribution. Based on the Wigner distribution function, we can further obtain the corresponding Wigner entropy and employ the feature that the critical state has the maximum Wigner entropy to locate the invariable mobility edge. Finally, we reveal that there are anomalous transport phenomena between the transition from ballistic transport to the absence of diffusion.
pacs:
71.23.An, 71.23.Ft, 05.70.Jk
I Introduction
Anderson localization Anderson (1958), a fundamental phenomenon in wave propagation through disordered media, continues to be an active field in condensed matter physics. Scaling theory Abrahams et al. (1979), applied to disordered systems, highlights the critical influence of spatial degrees of freedom on Anderson localization. In one- and two-dimensional systems, the introduction of uncorrelated random disturbances results in the exponential localization of all wave functions. Consequently, the localization-delocalization transition has been a longstanding focus in low-dimensional disordered systems. Conversely, three-dimensional (3D) Anderson system Anderson (1958); Mott (1987) displays a unique behavior, where wave functions are neither entirely localized nor delocalized. An energy threshold, known as mobility edge, separates delocalized states from localized ones. However, detecting the mobility edge in 3D systems experimentally remains difficult. Therefore, to gain a deeper understanding of the mobility edge, low-dimensional systems—particularly one-dimensional (1D) systems—offer a more feasible research path.
Except for the Anderson systems, the 1D Aubry-André (AA) model Aubry and André (1980) with a self duality plays a similarly important role in understanding the Anderson localization and the mobility edges as well. The AA model has an extended-localized transition point, which can be exactly derived by the dual transformation. In addition, AA model is also one of the important source for designing systems with mobility edges. The pursuit of systems with mobility edges has a rich history, including the development and study of numerous theoretical models over the years Abrahams et al. (1979); Mott (1987); Soukoulis and Economou (1982); Lee and Ramakrishnan (1985); Das Sarma et al. (1986, 1988, 1990); Liu et al. (2017a); Biddle and Das Sarma (2010); Biddle et al. (2011); Ganeshan et al. (2015); Liu and Guo (2018); Ye et al. (2024). Furthermore, recent experiments have successfully identified mobility edges in one-dimensional quasi-periodic optical lattices by manipulating cold atoms Lüschen et al. (2018); An et al. (2018); Kohlert et al. (2019). Specifically, Das Sarma et al. revealed that there was mobility edge in a class of systems with slow-varying quasidisordered on-site potentials Das Sarma et al. (1988, 1990); Liu et al. (2017a). Biddle et al. found the exactly solvable mobility edge in a generalized AA model with long-range tunnelings Biddle and Das Sarma (2010); Biddle et al. (2011). Ganeshan et al. verified the existence of the mobility edge protected by the self duality Ganeshan et al. (2015). Liu et al. obtained the exactly solvable mobility edge in the off-diagonal quasidisordered models Liu and Guo (2018). Wang et al. uncovered a class of exactly solvable 1D models with mobility edgesWang et al. (2020). Zhou et al. not only proposed a class of exactly solvable models which host a novel type of exact mobility edge, but also propose a novel experimental scheme to realize the exactly solvable model and the new mobility edges in an incommensurate Rydberg Raman superarray Zhou et al. (2023). Wang et al. obtained the analytical expression of the anomalous mobility edge, which separates localized states from critical states in the case with unbounded quasiperiodic potential Wang et al. (2023). Guan et al. studied localization properties and mobility edges of a generalized spinful Aubry-André-Harper (AAH) model, and analytically obtained exact expressions of the mobility edges by a semiclassical method.Guan et al. (2023). Liu investigated the properties of mobility edges in 1D incommensurate lattices with p-wave superfluidity Liu et al. (2022a). Ye et al. studied the mobility edge in a family of quasiperiodic systems evolving far from equilibrium Ye et al. (2024). Yamilov et al. investigated the Anderson transition for 3D light and identified a mobility edge that separates diffusive transport and Anderson localization Yamilov et al. (2025). Longhi demonstrated the existence of a mobility edge in quasicrystals induced by pure dephasing effects through the study of photonic quantum walks in synthetic mesh lattices Longhi (2024), and also provided an exact spectral analysis of the mosaic Wannier-Stark Hamiltonian, showing that there are no mobility edges strictly separating extended and localized states Longhi (2023). Qi et al. discovered quantum Griffiths singularities in ${\rm MgTi_{2}O_{4}}$, where the mobility edge of Anderson localization may lead to the occurrence of a 3D quantum phase transition and the emergence of quantum Griffiths singularities Qi et al. (2024). Chang et al. experimentally investigated a mosaic quasiperiodic photonic waveguide lattice and discovered the mobility edge Chang et al. (2025). Gonçalves et al. proposed a class of solvable one-dimensional quasi periodic tight binding models, in which the extended phase, localized phase, and critical phase are separated by mobility edges Gonçalves et al. (2023).
In addition to analytical demonstrations and experimental observations, the extended-localized transition can be observed by the dynamical evolution as well. Xu et al. have studied the wave packet dynamics of a one-dimensional system with generalized quasidisordered potential and parameter-dependent mobility edge Xu et al. (2020), and found that in the extended regime and the regime with mobility edge, the wave packet spreads with the behavior of ballistic transport, while in the localized regime, the wave packet does not diffuse. In fact, there is anomalous spreading phenomenon. It means that, as the change of systematic parameter, the spreading form of the wave packet may successively exhibit phenomena such as ballistic transport, superdiffusion, diffusion, subdiffusion, and non-diffusion Purkayastha et al. (2021). For example, for the Fibonacci model whose on-site potential is arranged with a Fibonacci substitution rule Chiaracane et al. (2021) and the AA model in isolated and open systems Purkayastha et al. (2018); Lacerda et al. (2021), there are anomalous transport phenomena. These findings drive us to inquire whether similar anomalous transport phenomena can occur in a system with a generalized quasidisordered potential and invariable extended-localized mobility edge. In 1932, Eugene Wigner preposed the phase space representation, i.e., Wigner distribution function $W(x,p)$ (here $x$ and $p$ denote the coordinate and momentum, respectively), which enables the concurrent description of quantum states or signals in both coordinate and momentum Wigner (1932); Wig (1984); Kenfack and Życzkowski (2004); Wig (2014); Simon (2000); Taghiabadi et al. (2016); Siyouri et al. (2016). It serves as a distinctive “fingerprint” that captures the joint distribution of these variables, hence providing a unique representation for quantum states or signals. This prompts us to further reveal the difference between the extended state and the localized state in both coordinate and momentum dimensions, and to further consider how to employ the Wigner distribution function to locate the mobility edge.
The rest of this paper are arranged as follows: In Sec. II we introduce the generalized AA model with a mobility edge and discuss the extended-localized transition. In Sec. III we introduce the Wigner distribution and discuss the correspondence between the Wigner distribution and the wave functions. Besides, we introduce the Wigner entropy and reveal that the mobility edge can be located by Wigner entropy. In Sec. IV, we study the quantum transport property of the model. Finally, we make a summary in Sec. V.
Figure 1: (Color Online) The eigenvalues $E$ of Eq. (2) versus $V$ under $b=0.4$, $L=500$. The color bar denotes the value of IPR. The blue solid line denotes the invariable mobility edge with $E_{c}=2t/b=5t$.
II Model and the mobility edge
We study a 1D generalized AA model with nearest-neighbor hoppings and incommensurate on-site potentials, whose Hamiltonian is expressed as
$\hat{H}=\sum^{L-1}_{n}t\left(\hat{c}^{{\dagger}}_{n+1}\hat{c}_{n}+\hat{c}^{{\dagger}}_{n}\hat{c}_{n+1}\right)+\sum^{L}_{n}V_{n}\hat{c}^{{\dagger}}_{n}\hat{c}_{n},$ (1)
where $\hat{c}^{{\dagger}}_{n}$ ($\hat{c}_{n}$) is the creation (annihilation) operator with $n$ the site index, $L$ is the length of the system and $t$ is the hopping strength which is set as the unit of energy. $V_{n}=\frac{V}{1-b\cos(2\pi\alpha n)}$ is the generally quasidisordered potential, in which $V$ is the strength of the potential with $10<b<$ 1 being the tuning parameter and $\alpha=\frac{\sqrt{5}-1}{2}$ is the incommensurate parameter. In coordinate space, we can define the generalized wave function |$\ket{\psi}=\sum_{n}\phi_{n}\hat{c}^{{\dagger}}_{n}\ket{0}$ with $\phi_{n}$ the amplitude of probability. Based on |$\ket{\psi}$, we can obtain the following static Schrödinger equation
$t\phi_{n+1}+t\phi_{n-1}+\frac{V}{1-b\cos(2\pi\alpha n)}\phi_{n}=E\phi_{n}.$ (2)
Employing the self-dual transformation, we can obtain the exact expression of the extended-localized mobility edge $E_{c}$ of the generalized AA model Liu et al. (2022b), which satisfying $E_{c}=\frac{2t}{b}$. From the exact expression of $E_{c}$, we know that the mobility edge is invariant with $V$, but is uniquely determined by $b$. In the following, we will show that in addition to the inverse participation ratio and finite-size scaling analysis to study the extended-localized transition over the energy domain, the extended-localized transition can be characterized by the Wigner distribution as well. Based on the Wigner distribution function, we can further define the winger entropy and show that the mobility edge can be located by winger entropy, and the results are consistent with the analytical solution. Finally, we will show that the system with generalized quasidisordered potential has the property of anomalous quantum transport.
Figure 2: (Color Online) (a)-(d) Wave functions obtained from Eq. (2) with $b=0.4$, $V=4.8t$ and $L=1000$. Concretely: (a) Extended state far below $E_{c}$; (b) and (c) critical states at $E_{c}$; (d) localized state far above $E_{c}$. (d) Finite-size scaling analysis on the fractal dimension $\beta_{\rm min}$ at $b=0.4$ and $V=4.8t$. $m$ is the index of the Fibonacci number.
Taking $b=0.4$ and $L=500$, the resulting localization phase diagram is plotted in Fig. 1. The color bar denotes the value of inverse participation ratio (IPR) defined as ${\rm IPR}_{j}=\sum^{L}_{n=1}|\phi^{j}_{n}|^{4}$ with $j$ the energy level index. For an extended state, ${\rm IPR}\propto L^{-1}$ and decays to zero under larger $L$. For a localized state, ${\rm IPR}$ is a finite value Biddle and Das Sarma (2010); Biddle et al. (2011). From Fig. 1, we can see that the invariable mobility edge $E_{c}/t=5$ perfectly separates the extended states from the localized state. Below $E_{c}$ ${\rm IPR}$ tends to 00, signaling the extended states. Above $E_{c}$, ${\rm IPR}$ are finite values, signaling the localized states. The numerical results are consistent with analytical results. To be specific, we plot the typically extended and localized state in Figs. 2(a) and 2(d), whose energy is far below $E_{c}$ and far above $E_{c}$, respectively. We can see that the probability distribution of the extended state extends over all the system while that of the localized state only occupies a small part of the system. In addition, at $E_{c}$, the wave functions are critical, because we can see from Figs. 2(b) and 2(c) that although the probability distribution occupies most of system space, there is still a part of the system space that is not occupied, presenting the characteristics of subexpansion.
[…]
Having recalled the extended-critical transition and the invariable caused by the generalized potentials, in the following section, we will focus on employing the Wigner distribution, a phase space pattern, to distinguish extended states and the localized states, and employing Wigner entropy to locate the mobility edge. In addition, we will reveal the anomalous transport phenomenon of the model.
Figure 3: (Color Online) The corresponding Wigner distributions of the wave functions under $b=0.4$, $V=4.8t$ and $L=1000$. (a) $W(x,p)$ of the $100$-th wave function; (b) $W(x,p)$ of the $900$-th wave function; (c) $W(x,p)$ of the $520$-th wave function; (d) $W(x,p)$ of the $527$-th wave function. The color bar represents the value of $W(x,p)$.
III Wigner distribution and Wigner entropy
For the given wave function |$\ket{\psi}$, the Wigner distribution function $W(x,p)$ can be obtained by the following integration Wigner (1932); Wig (1984); Kenfack and Życzkowski (2004); Wig (2014); Simon (2000); Taghiabadi et al. (2016); Siyouri et al. (2016)
[…]
where $x$ and $p$ are the coordinate and momentum in phase space, respectively, and $\hbar$ is the reduced Planck constant, and $\rho=|\psi\rangle\langle\psi|$.
The difference between the extended states and localized states can be easily seen from the Wigner distributions. We diagonalize the Hamiltonian matrix under $b=0.4$, $V=4.8t$, and $L=1000$. The $W(x,p)$ of an extended state (The $100$-th wave function) is plotted in Fig. 3(a). As seen that in the phase space, for the extended state, its $W(x,p)$ is extended and continuous in the $x$ branch, while in the $p$ branch, the distributions are discrete. For the localized state, the consequence is different. We take the $900$-th wave function (localized state) as an example, the corresponding $W(x,p)$ is plotted in Fig. 3(b). In the $x$ branch, the distributions are localized, but in the $p$ branch, the distributions are extended and continuous. In fact, for the critical states, the results are different from those of the extended and localized states as well. We take the $520$-th and $527$-th wave functions (critical states near $E_{c}$) as examples, whose corresponding $W(x,p)$ are plotted in Figs. 3(c) and 3(d), respectively. We can see that, the distributions are both subextended and segmented continuous in the $x$ and $p$ branches. We emphasis that the differences of $W(x,p)$ among extended, critical, and localized states are universal. For wave functions of other energy levels, the corresponding Wigner distribution results are similar to those in Fig. 3. Therefore, one can effectively employed the Wigner distribution to distinguish the extended, critical, and localized states.
Figure 4: (Color Online) (a)-(c) The Wigner entropy as the change of eigenenergies. (a) $V=3.2t$; (b) $V=4.8t$; (c) $V=7.3t$. (d) $W_{S}$ in the $E$-$V$ parameter space with solid blue line the $E_{c}=5t$. Other involved parameters are $b=0.4$ and $L=1000$.
Based on the Wigner distribution, one can further obtain the Wigner entropy (marked by $W_{S}$) Van Herstraeten and Cerf (2021). Considering the negativity of $W(x,p)$, the $W_{S}$ is finally calculated by the following definition
$W_{S}=-\iint W(x,p)\ln{|W(x,p)|}dxdp,$ (4)
where the integral region of $x$ is $\left[-L/2,L/2\right]$, and that of $p$ is $\left[-\pi,\pi\right]$. $W_{S}$ can reflect whether the distribution of quantum states in phase space is concentrated or not. If Wigner entropy is small, the distribution of quantum states is concentrated. If $W_{S}$ is large, it means that the distribution of quantum states is relatively dispersed. We notice that the $W(x,p)$ of the localized and extended states is more concentrated than that of the critical states (As seen from Fig. 3, the $W(x,p)$ corresponding to the localized state and the extended state are relatively finite and concentrated on the position and momentum branches, respectively, while the distributions of critical sates are relatively dispersed.), so we infer that the quantum state has the largest $W_{S}$ at the mobility edge. In addition, we can see from Figs. 3(a) and 3(b) that the fringe number of $W(x,p)$ of the localized state is obviously smaller than that of the extended state, therefore, we infer that the localized state has the smallest $W_{S}$. According to our inference, It seems possible to locate the mobility edge by $W_{S}$.
To verify our inference, we plotted $W_{S}$ for $V=3.2t$, $V=4.8t$, and $V=7.3t$ in Figs. 4(a), 4(b), and 4(c), respectively. From the figures, we can see that $W_{S}$ peaks at $E_{c}=5t$ which means that the critical state has a maximal $W_{S}$ at $E_{c}$. Moreover, the $W_{S}$ below $E_{c}$ is slightly larger than that above $E_{c}$. Concretely, the maximal $W_{S}$ for the $V=4.8t$ case is $W^{critical}_{S}=5.2217$. Besides, we make averages on $W_{S}$ before and after $E_{c}$, respectively, and obtain the averaged $W_{S}$ for the extended states satisfies $W^{extended}_{S}=3.7875$ and that for the localized states is $W^{localized}_{S}=2.6249$. From the numerical results, it is intuitive that $W^{critical}_{S}>W^{extended}_{S}>W^{localized}_{S}$. Therefore, the inference made from the distributions of $W(x,p)$ is correct under the validations of numerical calculation. We emphasize that this conclusion is universal to the results for other $V$ and $b$. In addition, the results tell that we can locate the mobility edge by using the property that the mobility edge produces the maximum Wigner entropy, and distinguish the extended state from the local state by comparing the average entropy of the left and right sides of the maximum. Meanwhile, a full phase diagram that contains $W_{S}$ in the $E$-$V$ parameter space has been presented in Fig. 4(d). The color bar represents $W_{S}$, and from the color point of view, the mobility edge can separate the two states of extension and localization. Recently, there are different ways to realize the measurement the Wigner distribution function, such as the qubit quantum processor Tilma et al. (2016); Rundle et al. (2017), discrete atomic system Tian et al. (2018), photonic system Tilma et al. (2022); Xu et al. (2014), and optical parametric amplification Kalash and Chekhova (2023). Wigner entropy is a statistical result of Wigner distribution. With these experimental techniques, it may be possible to experimentally locate the mobility edge accurately in the future.
Figure 5: (Color Online) The time evolution of mean squared displacement $\Delta x^{2}(\tau)$ under $b=0.4$ and $L=1000$. (a) $V=0.2t$; (b) $V=2.5t$; (c) $V=4.8t$; (d) $V=6.8t$; (e) $V=7.3t$; (f) $V=10t$.
IV Quantum transport property
In the previous study, it was proposed that the mobility edge can observed by the dynamical observation Xu et al. (2020) in an one-dimensional incommensurate system with generalized quasidisordered potential, where extended-critical transition can be characterized by the transition from the ballistic transport to the absence of diffusion. In fact, in addition to the ballistic transport and absence of diffusion, there are anomalous transport phenomena in the incommensurate system Purkayastha et al. (2021); Chiaracane et al. (2021); Purkayastha et al. (2018); Lacerda et al. (2021), such as the superdiffusive and subdiffusive transports. Therefore, in this section, we are concerned about whether anomalous transport occurs in the system with the generalized quasidisordered potential similar to that in Ref. Xu et al. (2020) and with the invariable mobility edge.
We analyze the transport behaviors by studying the dynamical evolution of the wave packet. The isolated generalized AA model is initialized with a particle occupying on site $L/2$. Thus, the initial state |$\ket{\psi(\tau=0)}$ is described by the wave function |$\ket{\psi(\tau=0)}=\sum_{n}\phi_{n}(\tau=0)\hat{c}^{{\dagger}}_{n}\ket{0}$ where $\phi_{n}(\tau=0)=\delta_{n,L/2}$ is the Kronecker $\delta$ function. By solving the time-dependent Schrödinger equation, we obtain the evolution of the wave function probability amplitude with time $\phi_{n}(\tau)$, and then the mean squared displacement of wave function is given by Chiaracane et al. (2021)
$\Delta x^{2}(\tau)=\sum^{L}_{n=1}\left(n-L/2\right)^{2}|\phi_{n}(\tau)|^{2}.$ (5)
The asymptotic relation of square mean displacement can also be written as
$\Delta x^{2}(\tau)\sim\tau^{\gamma},$ (6)
where $\gamma=2$ represents the ballistic transition, $\gamma=0$ implies absence of diffusion, $\gamma=1$ denotes diffusion, $10<\gamma<$ 1 shows subdiffusive transport, and $21<\gamma<$ 2 corresponds to superdiffusive transport. In Figs. 5(a)-5(f), we plot the the time evolution of $\Delta x^{2}(\tau)$ under $b=0.4$ and $L=1000$. From the results of $\Delta x^{2}(\tau)$, we conclude the in this one-dimensional incommensurate system with generalized quasidisordered potential and invariant mobility edge, the evolution of wave packet is not only ballistic transport and absence of diffusion, but also anamolous transport phenomena. Specifically, when the potential strength is small, such as $V=0.2t$, the system is in the extended phase. Then we can see from Fig. 5(a) that the evolution of wave packet presents ballistic transport. As $V$ get larger, the system entries the mobility edge regime, there are anomalous transport phenomena. From Figs. 5(b)-5(e), we can see that the evolution curves of $\Delta x^{2}(\tau)$ presents superdiffusion, diffusion, and subdiffusion transport in turn. In the localization regime, such as $V=10t$, $\Delta x^{2}(\tau)$ scales as $\tau^{0}$ (see Fig. 5(f)), meaning the absence of diffusion.
V Summary
In conclusion, we have studied a generalized Aubry-André model with tunable quasidisordered potentials. The model exists an invariable mobility edge that can separate the extended state from the localized state. At the invariable mobility edge, the wave functions are critical. We have proved that one can employ the physical image in phase space, i.e., the Wigner distribution, to distinguish the extended states, critical states, and the localized states. Based on the Wigner distribution, we can further obtain the Wigner entropy. Our numerical results show that it is also feasible to use Wigner entropy to distinguish extended, critical and local states. The critical state has the largest entropy, the extended state the middle, and the local state the smallest. Using this property, we can locate the invariable mobility edge. In addition, by studying the wave packet dynamics of this model, we find that there are not only ballistic transport, diffusion, and absence of diffusion, but also are superdiffusion and subduffusion. These results further expand the extended-localized transition and dynamic properties of generalized quasidisordered systems. In the near future, we hope that these approaches can be generalized to other disordered or quasidisordered systems to distinguish between different forms of phase transitions.
Acknowledgements.
This work is supported by the Natural Science Foundation of Zhejiang Province (Grants No. LQN25A040012) and the start-up fund from Xingzhi college, Zhejiang Normal University, and the National Natural Science Foundation of China (Grants No. 12174346),